Assuming that a gas X is an ideal gas. If 22.4g of X occupies 11.2dm at stp
a)How many moles of X are in the system?
b) what is the molecular weight of X
c)how many molecules of X are in the system?
d)What will be the volume occupied by the gas if it's heated to a temp of 273°c and the pressure is simultaneously reduced to half of the original value
1
Expert's answer
2020-04-22T09:17:09-0400

22.4 g gas occupies 11.2 dm .It means

44.8 g gas will occupy 22.4 dm at STP. So molar mass = 44.8 gm per mole

(a) moles of X ="\frac{mass}{molar \ mass}=\frac{22.4}{44.8}= 0.5 \ moles"


(b) molecular weight = molar mass = 44.8 gm per mole


(c)from ideal gas equation


"\frac{PV}{T}= constant"

So,


"P_1V_1T_2=P_2V_2T_1\\1\times11.2\times546K=0.5\times V_2\times 273\\or\\V_2=44.8\ dm"